OpenStudy (anonymous):

p(x)=x^3-4x^2+3x+1, q(x)=x^2-2x-4 Find the coordinates of the intersection point of the two curves in QIII. What does this point represent with respect to the equations?

OpenStudy (anonymous):

I think if someone told me how to start it I might be able to finish, but I'm not really sure.

OpenStudy (anonymous):

To find points of intersection just set p(x) = g(x) and solve for x. The point will represent where the two graphs cross each other.

OpenStudy (anonymous):

\[x^3-4x^2+3x+1 = x^2-2x-4 \] \[\rightarrow x^3 - 5x^2 + 5x + 5 = 0\] Factor this and we get \[\rightarrow x=−0.598675\] Plug this into p(x) or g(x) to get y -value and that is your point of intersection.

OpenStudy (anonymous):

Thank you so much @arjund12 That was really helpful.